OpenStudy (anonymous):

Does every nth degree polynomial have (n-1) critical numbers? Why or why not?

OpenStudy (anonymous):

nth degree polynomial has the form Pn(x)=a0+a1x+......+a(subscript n)x^n coefficient can differ in number from 1 to n... with a(subscript n) never equal to zero

OpenStudy (anonymous):

so, why ids it false?

OpenStudy (anonymous):

*is

OpenStudy (sirm3d):

a big NO! \[\large y=x^3+3x\]has no critical number

OpenStudy (anonymous):

can u give me another example besides that

OpenStudy (sirm3d):

\[\large y =x^5 + 5x\]

OpenStudy (anonymous):

stgreen, is that true?

OpenStudy (sirm3d):

why don't you test it yourself. take the derivative of the function, then try to find the real roots or critical numbers.

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

ur right!

OpenStudy (anonymous):

but can u give me a really good explanation on why its false?

OpenStudy (anonymous):

nvrm.......how bout this: does an nth degree polynomial has at most (n-1) critical numbers

OpenStudy (sirm3d):

some derivatives, particularly quadratic polynomials, have no real roots, and we say the function has no critical number.

OpenStudy (anonymous):

okay thnx guys

OpenStudy (sirm3d):

at most (n-1) critical numbers allow 0, 1, 2, 3, .. up to (n - 1) critical numbers. This is certainly true.

OpenStudy (anonymous):

ya...i got that one